OpenStudy (anonymous):

Help? Pineapples cost $2.50 each and mangos cost $1.50 each. Tim purchased 15 pieces of fruit for $29.50. How many pineapples and how many mangos did he purchase?

OpenStudy (jhannybean):

let x = pineapples, and y = mangoes. first equation: states you have # of pineapples + # of mangoes = total # of fruit. Second equation: states cost of pineapples + cost of mangoes = total cost for both pineapples and mangoes.

OpenStudy (anonymous):

is this equation right? 15 x 2.50x + 1.50y

OpenStudy (anonymous):

@Jhannybean

Directrix (directrix):

Is there supposed to be an equality symbol in here? Just wondering about that. 15 x 2.50x + 1.50y

OpenStudy (anonymous):

i was going to put one but i didnt know what to put after it

OpenStudy (jhannybean):

Just seems like a system of equation problem to me.

Directrix (directrix):

I started out this way: Let p = the number of pineapples 15- p is the number of mangos So, 2.50*p + 1.50*(15 - p) = 29.50

OpenStudy (anonymous):

I will try that

OpenStudy (jhannybean):

I was solving it this way: let x = pineapples, and y = mangoes. x+ y = 15 2.50x + 1.50y = 29.50

OpenStudy (anonymous):

I dont really get how i figure out how many pineapples and how many mangoes there are?

OpenStudy (jhannybean):

You solve it like a system of linear equations by the process of elimination.

OpenStudy (jhannybean):

Or substitution.

OpenStudy (jhannybean):

ou want to eliminate 1 variable in order to solve by the process of elimination, so we could start by getting rid of the x's first. Multiply by the reverse value. -2.50(x+ y = 15) 2.50x + 1.50y = 29.50 _________________________ -2.50x - 2.50y = -37.50 2.50x + 1.50y = 29.50 _______________________________ 0x - 1y = -8 y = 8 mangoes

OpenStudy (anonymous):

ok since theres 8 mangoes that will mean theres 7 pineapples? right?

OpenStudy (jhannybean):

Now you can use your y-value into the first equation to solve for number of pineapples. Yes, 7 pineapples :) \(\checkmark\) x + 8 = 15 x = 15 - 8 = 7 pineapples Good job.

OpenStudy (anonymous):

:) thank you so much! i was stuck on this problem forever

OpenStudy (jhannybean):

Haha, no problem!

OpenStudy (jhannybean):

Yes, @Directrix's method works as well.